OpenStudy (anonymous):

How do I find the domain and range of 2x^2-7?

OpenStudy (anonymous):

Hey there. That equation bases from the original formula of \[y=ax ^{2}+bx+c\]., where a is the value attached to the x^2 variable, b is the value attached to the x variable and c is attached to no variable. Often, the domain of a parabola is \[x\], which also applies to this one. To find y-vertex, which is the tip of the parabola is by using the formula\[y _{v}=\frac{ b ^{2}-4ac }{ 4a }\]. After this, we replace the values, being\[y _{v}=\frac{ 0^{2}-4*2*(-7) }{ 4*2 }=7\]. It is also important to take account of the a value. When it is positive, the parabola is U-shaped and when it is negative, it is cap-shaped. In this case the a is positive, giving us the clue that the vertex is at the bottom, with y-value equivalent to 7. It then goes up to infinity. Hence, the range is \[y \ge7\]

OpenStudy (anonymous):

The parabola of the graph is \[x \in \mathbb{R} \]

OpenStudy (anonymous):

That is the domain.